Write an equation for the set of all points in the plane equidistant from (0, – 1) and y= – 5. Write your answer in vertex form. Simplify any fractions
1
Expert's answer
2020-04-27T17:40:45-0400

Let a point P = "(x, y)" .

Given,

A = (0, -1) and

Equation of a line is "\space y + 5 =0"


PA = "\sqrt {(x_2-x_1)^2 + (y_2 -y_1)^2} = \sqrt {(0-x)^2 + (-1-y)^2}"


PA = "\sqrt {x^2 + y^2 + 2y+ 1}"


Perpendicular Distance from the point "(x,y)" to the Line L: "y +5 = 0"


PL = "|\frac{x(0) +y(1) +5}{\sqrt {0+1}} | = |{y+5}|"


Given, PA = PL



"PA^2 = PL^2"

"x^2+y^2 +2y +1 = (y+5)^2"

"x^2 +y^2 + 2y +1 = y^2 +10y + 25"

"x^2 = 8y+24"

"x^2 = 8 (y +3)"

The equation of set of points in vertex form is, "x^2 = 8(y+3)"